OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (1, –3) and has a slope of 2?

OpenStudy (zpupster):

y=mx+b m= slope so y= 2x+b y-(-3) = 2(x-1) or y+3 = 2x-2 or y= 2x-5